If a garden is 20 yd by 16 yd, what are it's dimension in meters?

With solution please...
i need the answer after 2 hours. i really really need it..
Thank you in advance

Answer by nerdybill(7384)

20 yd by 16 yd
.
Since:
3 feet per yard
and
1 meter per 3.2808399 feet
.
20 yd = (20*3)/3.2808399 = 18.29 meters
16 yd = (16*3)/3.2808399 = 14.63 meters
.
18.29 meters by 14.63 meters